The owner of a shop manufactures two products: Product B (B) and Product C (C). Each Product B requires 6 ounces of ingredient I, 1 gram of ingredient II and 2 tablespoons of ingredient III. A Product C requires 3 ounces of ingredient I, 1 gram of ingredient II and 4 tablespoons of ingredient III. The company has 6600 ounces of ingredient I, 1400 grams of ingredient II and 4800 tablespoons of ingredient III. Profits for Product B are $0.20 per unit, and for Product C $0.30 per unit. If the owner manufactures 600 units of Product B and 800 units of Product C, which resource has “slack” (i.e., which resource is not fully used)?

A. Ingredient I (only)

B. Ingredient I and ingredient II

C. Ingredient III and ingredient I

D. Ingredient III (only)

E. Ingredient III and ingredient II
Answer #1

C. Ingredient III and ingredient I

Product B

Total Units produced = 600 Units

Ingredient 1 = 6*600

= 3600 Ounces

Ingredient II = 1*600

= 600 Grams

Ingredient III = 2*600

= 1200 Tablespoons

Product C

Total Units produced = 800

Ingredient I = 3*800

= 2400 Ounces

Ingredient II = 1*800

= 800

Ingredient III = 4*800

= 3200

Hence it can be observed that based on the situation, ingredient I and III are not completely used for the same.